OpenStudy (blaque678):

The perimeter of the rectangle is 98 units.Find the value of x. 5x+1 4x+3.

terenzreignz (terenzreignz):

Do 5x + 1 and 4x + 3 refer to the length and width of the rectangle?

OpenStudy (blaque678):

yes

terenzreignz (terenzreignz):

Well, do you know your formula for perimeter? :)

OpenStudy (blaque678):

2L+2W i believe

terenzreignz (terenzreignz):

That's right :) Now if we just let L = 5x + 1 and W = 4x + 3 What does that give us? :D

terenzreignz (terenzreignz):

(by the way) if you let L = 4x +3 and W = 5x + 1 it really doesn't change anything :)

OpenStudy (blaque678):

I get mixed up on the math part thats why i asked for help!!!!

terenzreignz (terenzreignz):

Oh ok :) As we know our length is L = 5x + 1 and our width is W = 4x + 3 We know that the formula for perimeter is P = 2L + 2W We also know that the perimeter is 98. So we just replace P with 98 L with 5x + 1 and W with 4x + 3 Which brings us to: 98 = 2(5x + 1) + 2(4x + 3) Can you do the rest? :)